6. A 40.0-W lightbulb is connected to a 220.0-V source.

(a) What is the current through the bulb?
A

(b) What is the internal resistance of the bulb?

?

If you could provide a step by step solution, that would be great.
Answer #1

Ans: Given Power consumed by the bulb of resistance R is P= 40 W

Voltage applied is V= 220 V

a) Current through the bulb is I= P/V = 40/220 = 0.1818 A = 0.182 A= 182 mA

b) Internal resistance of the bulb is R= V/I = 220/ 0.182 = 1210 ohms